OpenStudy (anonymous):

Simplify a. i^23 b. i^42

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i^23 = i.i^22 = i(i^2)^11 = do the rest i^42 = (i^2)^21 = do it

OpenStudy (anonymous):

I don't get it...?

OpenStudy (anonymous):

ok let me explain as i = sqrt(-1) and i^2 = -1 now i^23 = i x i^22 = i x (i^2)^11 = i x (-1)^11 = do the last step now

OpenStudy (anonymous):

now for the second question i^42 = (i^2)^21 = (-1)^21 simplify it get your answer got it??????

OpenStudy (anonymous):

so for the first one it would be -1i? @rizwan_uet

OpenStudy (anonymous):

and for the second one it would be -1? @rizwan_uet

OpenStudy (anonymous):

yes you are right a) -i b) -1

OpenStudy (anonymous):

can you explain this in words step by step please? So I can remember.

OpenStudy (anonymous):

and isn't -i simply just -1?

OpenStudy (anonymous):

no no -i not equals -1

OpenStudy (anonymous):

oh okay nevermind then.

OpenStudy (anonymous):

do you know i (iota) ???

OpenStudy (anonymous):

nope...